Given f(x) = 3x + 3, find f(a + 1).
A) a + 7 B) a + 6 C) 3a + 4 D) 3a + 6

Answer by scott8148(6628)

f(a+1)=3(a+1)+3 ... f(a+1)=3a+3+3 ... f(a+1)=3a+6